# PhysioWave: A Multi-Scale Wavelet-Transformer for Physiological Signal Representation

### v2

@@ Abstract
Physiological signals are often corrupted by motion artifacts, baseline drift, and other low-SNR disturbances, which pose significant challenges for analysis. Additionally, these signals exhibit strong non-stationarity, with sharp peaks and abrupt changes that evolve continuously, making them difficult to represent using traditional time-domain or filtering methods. To address these issues, a novel wavelet-based approach for physiological signal analysis is presented, aiming to capture multi-scale time-frequency features in various physiological signals. Leveraging this technique, two large-scale pretrained models specific to EMG and ECG are introduced for the first time, achieving superior performance and setting new baselines in downstream tasks. Additionally, a unified multi-modal framework is constructed by integrating pretrained EEG model, where each modality is guided through its dedicated branch and fused via learnable weighted fusion. This design effectively addresses challenges such as low signal-to-noise ratio, high inter-subject variability, and device mismatch, outperforming existing methods on multi-modal tasks. The proposed wavelet-based architecture lays a solid foundation for analysis of diverse physiological signals, while the multi-modal design points to next-generation physiological signal processing with potential impact on wearable health monitoring, clinical diagnostics, and broader biomedical applications.
@@ 1 Introduction
Multi-modal biosignal learning faces two key challenges: first, extreme heterogeneity across modalities like EEG, EMG, and ECG, which operate at different sampling rates and temporal scales, leading to potential misalignment and spectral aliasing [21]; second, variability in signal quality due to motion artifacts or electrode drift [22]. To overcome these, we use modality-specific backbones: a pretrained encoder for EEG and newly pretrained PhysioWave encoders for EMG and ECG. These backbones remain frozen during downstream training, with only lightweight classification heads and fusion coefficients being learned. The fusion layer dynamically adjusts weights to prioritize reliable modalities, ensuring robust predictions. This linear-probing approach consistently outperforms single-modality baselines across multi-modal tasks (see Figures 5 and 6), including a 7.3% gain in classification accuracy on DEAP [23], showcasing the effectiveness of dynamic, reliability-aware fusion for heterogeneous physiological data.
@@ 2 Method
Model Architecture: Figure 1 illustrates the end-to-end model pretraining pipeline. The raw multi-channel signal is first segmented into overlapping windows to form samples $x\in\mathbb{R}^{C\times T}$, where $C$ denotes the number of channels and $T$ the number of time steps. Each sample undergoes a learnable wavelet decomposition across $L$ levels, yielding multi-scale frequency-band representations: detail subbands $d^{(l)}[n]\in\mathbb{R}^{C\times T}$ for $l=1,\dots,L$ and a final approximation subband $a^{(L)}[n]\in\mathbb{R}^{C\times T}$. Concatenation of these $L+1$ subband outputs along the channel axis produces a feature map $\mathrm{Spec}(X)\in\mathbb{R}^{((L+1)\,C)\times T}$ (Section  2.1). Next, the feature map is partitioned into uniform spatio-temporal patches: for each patch we compute its FFT-based spectral energy, blend this value with random noise to produce an importance score, and mask the highest-scoring patches by replacing their embeddings with a learnable <MASK> token. The remaining patches are projected into token embeddings and augmented with rotary positional embeddings (Section  2.2). Finally, the complete token sequence is passed through a stack of Transformer encoder layers, and a lightweight decoder reconstructs the masked patches (Section  2.3). In addition, end-to-end fine-tuning is performed downstream for single-modality tasks, while in multimodal settings modality-specific predictions are aggregated to yield the final output (Section  2.4).
@@ 2.1 Learnable Wavelet Decomposition
As depicted in the top-left corner of Figure 1, the model maintains $M$ candidate wavelet bases $\{(k_{w}^{\mathrm{low}},k_{w}^{\mathrm{high}})\}_{w=1}^{M}$ to accommodate diverse signal characteristics. For each input $x\in\mathbb{R}^{C\times T}$, temporal information is aggregated via average pooling and the resulting feature vector is passed through a compact MLP to produce unnormalized selection scores. Applying a softmax yields selection weights
@@
These aggregated filters serve as the effective low- and high-pass filters for all subsequent analysis and downsampling operations, and during training their weights continue to be updated to adapt to the actual signal distribution.
@@
At each decomposition level $\ell=1,\dots,L$, the gated approximation–detail pair $[\hat{a}^{(\ell)}[n],\hat{d}^{(\ell)}[n]]$ is concatenated along the channel axis and reshaped into a $\mathbb{R}^{2C\times 1\times T}$ feature map. This map is fed to a Cross-Scale Channel-Aggregation Feed-Forward Network (CAFFN; see Fig. 3[26]. CAFFN first applies a lightweight channel-aggregation block and then performs multi-head attention where the current features act as queries and the flattened feature maps from all shallower levels provide keys and values. Formally, let $\mathbf{U}^{(\ell)}$ be the CAFFN output before cross-scale fusion; the refined representation is obtained as
@@
where the learnable scalar $\beta$ balances the current-level information with the context aggregated from coarser resolutions. This cross-scale fusion enables fine-grained subband features to be informed by long-range patterns captured at earlier levels, yielding scale-aware, frequency-aware representations for subsequent stages.
@@ 2.2 Frequency-guided Masking and Tokenisation
2:$N\leftarrow\lfloor T/w\rfloor$triangleright number of patches
@@
5:for $n=1$ to $N$ do $\triangleright$ score each patch
6:     $F\leftarrow\operatorname{FFT}(p_{n})$triangleright frequency spectrum
7:     $e_{n}\leftarrow\sum_{k}\lvert F_{k}\rvert$triangleright spectral energy
@@
11:$s_{n}\leftarrow\alpha\,e_{n}+(1-\alpha)\,z_{n}$triangleright blended score
12:$\mathbf{i}\leftarrow\operatorname{argsort}(\mathbf{s})$triangleright ascending: low score first
13:$L\leftarrow\lfloor(1-\rho)N\rfloor$triangleright patches to keep
@@
15:Initialise $\mathbf{m}\leftarrow\mathbf{1}_{N}$; $\mathbf{m}[\text{mask}]\leftarrow 0$triangleright binary mask
@@ 2.3 Reconstruction
Patch-level reconstruction loss.
Let $\{\mathbf{p}_{n}\}_{n=1}^{N}$ and $\{\hat{\mathbf{p}}_{n}\}_{n=1}^{N}$ be the ground-truth and reconstructed frequency-band patches defined in Section 2.2, and let =$\mathcal{M}=\{n\mid m_{n}=1\}$ be the masked-patch index set. The model is trained to minimise the mean Smooth-L1 discrepancy over only the masked patches  [33, 34]:
@@
This objective 14 compels the network to recover the multiscale, frequency-band information concealed by the mask.
@@ 3.1 Datasets and Training Settings
Leveraging the PhysioWave architecture, two large-scale, modality-specific foundation models, PhysioWave-ecg and PhysioWave-emg, are pretrained on the most extensive open-access corpora currently available for their respective signal types (see Tables 7 and 8). PhysioWave-ecg is trained on approximately 182 GB of twelve-lead ECG recordings, while PhysioWave-emg utilizes about 823 GB of EMG data. For each modality, we provide three parameter configurations: Small (5M), Base (15M), and Large (37M). Both PhysioWave models share the same Transformer backbone; however, their learnable wavelet front-ends are modality-aware (see Tables 4 and 5 in Appendix B for full architectural and training details).
@@
Each recording is denoised with a modality-specific band-pass filter followed by a 50 Hz notch. Traces with fewer than 12 ECG leads or 16 EMG electrodes are zero-padded and then resampled to 500 Hz (ECG) or 2 kHz (EMG). See Appendix C and D for full details.
@@ 3.2 Downstream Experiment Results
Table 1: ECG rhythm classification.
Method (year) Params PTB-XL CPSC 2018 Chapman–Shaoxing F1 AUROC F1 AUROC F1 AUROC ECG-Chat (2024) 13 B 55.9 94.1 80.1 95.7 MERL [36] (2024) 11 M 48.1 91.9 72.8 92.6 87.9 MaeFE (2023) [37] 9 M 64.7 88.6 71.6 94.5 OpenECG-SimCLR 11 M 46.9 91.5 73.1 92.4 52.3 95.1 OpenECG-BYOL 11 M 47.7 91.1 72.8 92.6 51.5 94.8 OpenECG-MAE 11 M 48.1 90.9 74.5 93.2 50.8 94.2 Ours–Small 5 M 65.8 92.7 71.6 95.5 52.1 96.4 Ours–Base 15 M 64.5 93.4 72.5 95.9 53.8 97.2 Ours–Large 37 M 66.7 94.6 73.1 96.1 54.8 98.3
EMG Classification Results.
Table 2: Surface-EMG gesture recognition.
Method (year) Params NinaPro DB5 EPN-612 UCI EMG Acc. F1 Acc. F1 Acc. F1 Moment (2024) 385 M 86.41 74.42 90.87 90.16 90.45 91.75 OTiS (2024) 45 M 85.31 72.61 87.55 88.03 90.62 89.28 Ours–Small 5 M 84.78 72.54 93.12 93.40 90.35 89.51 Ours–Base 15 M 86.02 73.78 93.68 93.91 91.92 92.77 Ours–Large 37 M 87.53 75.42 94.50 94.56 93.19 93.59
@@
On MPDB, the same multimodal architecture—obtained by replacing the EEG branch with LaBraM ($5.8$ M parameters)—improves accuracy from $70.4\%$ (using LaBraM alone) to $74.9\%$ ($+4.5\%$), as shown in Figure 6. This outperforms generic sequence models such as MMPNet and EEGNet by $9\%$ to $17\%$  [39, 41].
Figure 5: Results on the DEAP dataset.
Figure 6: Results on the MPDB dataset.
@@ 5 Acknowledgment
This research was partly supported by the EU Horizon Europe project IntelliMan (g.a. 101070136), the PNRR MUR project ECS00000033ECOSISTER, and the ETH Zürich’s Future Computing Laboratory funded by a donation from Huawei Technologies. Furthermore, the research was also partially supported by the EU Horizon Europe project HAL4SDV (g.a. 101139789).
Appendix A RELATED WORK
A.1 Self-supervised pre-training for Time Series

### v4

@@ Abstract
Physiological signals are often corrupted by motion artifacts, baseline drift, and other low-SNR disturbances, which pose significant challenges for analysis. Additionally, these signals exhibit strong non-stationarity, with sharp peaks and abrupt changes that evolve continuously, making them difficult to represent using traditional time-domain or filtering methods. To address these issues, a novel wavelet-based approach for physiological signal analysis is presented, aiming to capture multi-scale time-frequency features in various physiological signals. Leveraging this technique, two large-scale pretrained models specific to EMG and ECG are introduced for the first time, achieving superior performance and setting new baselines in downstream tasks. Additionally, a unified multi-modal framework is constructed by integrating pretrained EEG model, where each modality is guided through its dedicated branch and fused via learnable weighted fusion. This design effectively addresses challenges such as low signal-to-noise ratio, high inter-subject variability, and device mismatch, outperforming existing methods on multi-modal tasks. The proposed wavelet-based architecture lays a solid foundation for analysis of diverse physiological signals, while the multi-modal design points to next-generation physiological signal processing with potential impact on wearable health monitoring, clinical diagnostics, and broader biomedical applications.  Code and data are available at: github.com/ForeverBlue816/PhysioWave
@@ 1 Introduction
Multi-modal biosignal learning faces two key challenges: first, extreme heterogeneity across modalities like EEG, EMG, and ECG, which operate at different sampling rates and temporal scales, leading to potential misalignment and spectral aliasing [21]; second, variability in signal quality due to motion artifacts or electrode drift [22]. To overcome these, we use modality-specific backbones: a pretrained encoder for EEG and newly pretrained PhysioWave encoders for EMG and ECG. These backbones remain frozen during downstream training, with only lightweight classification heads and fusion coefficients being learned. The fusion layer dynamically adjusts weights to prioritize reliable modalities, ensuring robust predictions. This linear-probing approach consistently outperforms single-modality baselines across multi-modal tasks (see Figure 5, including a 7.3% gain in classification accuracy on DEAP [23], showcasing the effectiveness of dynamic, reliability-aware fusion for heterogeneous physiological data.
@@ 2 Method
Model Architecture: Figure 1 illustrates the end‐to‐end model pretraining pipeline. The raw multi‐channel signal is first segmented into overlapping windows to form samples $x\in\mathbb{R}^{C\times T}$, where $C$ denotes the number of channels and $T$ the number of time steps. Each sample undergoes a learnable wavelet decomposition across $L$ levels, yielding multi‐scale frequency‐band representations: detail subbands $d^{(l)}[n]\in\mathbb{R}^{C\times T}$ for $l=1,\dots,L$ and a final approximation subband $a^{(L)}[n]\in\mathbb{R}^{C\times T}$. Concatenation of these $L+1$ subband outputs along the channel axis produces a feature map $\mathrm{Spec}(X)\in\mathbb{R}^{((L+1)\,C)\times T}$ (Section  2.1). Next, the feature map is partitioned into uniform spatio-temporal patches: for each patch we compute its FFT-based spectral energy, blend this value with random noise to produce an importance score, and mask the highest-scoring patches by replacing their embeddings with a learnable <MASK> token. The remaining patches are projected into token embeddings and augmented with rotary positional embeddings (Section  2.2). Finally, the complete token sequence is passed through a stack of Transformer encoder layers, and a lightweight decoder reconstructs the masked patches (Section  2.3). In addition, end-to-end fine-tuning is performed downstream for single-modality tasks, while in multimodal settings modality‐specific predictions are aggregated to yield the final output (Section  2.4).
@@ 2.1 Learnable Wavelet Decomposition
As depicted in the top‐left corner of Figure 1, the model maintains $M$ candidate wavelet bases $\{(k_{w}^{\mathrm{low}},k_{w}^{\mathrm{high}})\}_{w=1}^{M}$ to accommodate diverse signal characteristics. For each input $x\in\mathbb{R}^{C\times T}$, temporal information is aggregated via average pooling and the resulting feature vector is passed through a compact MLP to produce unnormalized selection scores. Applying a softmax yields selection weights
@@
These aggregated filters serve as the effective low‐ and high‐pass filters for all subsequent analysis and downsampling operations, and during training their weights continue to be updated to adapt to the actual signal distribution.
@@
At each decomposition level $\ell=1,\dots,L$, the gated approximation–detail pair $\hat{a}^{(\ell)}[n],\hat{d}^{(\ell)}[n]]$ is concatenated along the channel axis and reshaped into a $\mathbb{R}^{2C\times 1\times T}$ feature map. This map is fed to a Cross-Scale Channel-Aggregation Feed-Forward Network (CAFFN; see Fig. 3[26]. CAFFN first applies a lightweight channel-aggregation block and then performs multi-head attention where the current features act as queries and the flattened feature maps from all shallower levels provide keys and values. Formally, let $\mathbf{U}^{(\ell)}$ be the CAFFN output before cross-scale fusion; the refined representation is obtained as
@@
where the learnable scalar $\beta$ balances the current‐level information with the context aggregated from coarser resolutions. This cross-scale fusion enables fine-grained subband features to be informed by long-range patterns captured at earlier levels, yielding scale-aware, frequency-aware representations for subsequent stages.
@@ 2.2 Frequency-guided Masking and Tokenisation
[…]
@@
5:for $n=1$ to NN do $\triangleright$ score each patch
6:  $F\leftarrow\operatorname{FFT}(p_{n})$ $\triangleright$ frequency spectrum
7:  $e_{n}\leftarrow\sum_{k}\lvert F_{k}\rvert \triangleright$ spectral energy
@@
[…]
15:Initialise $\mathbf{m}\leftarrow\mathbf{1}_{N}$; $\mathbf{m}[\text{mask}]\leftarrow 0$ $\triangleright$ binary mask
@@ 2.3 Reconstruction
Patch‑level reconstruction loss.
Let $\{\mathbf{p}_{n}\}_{n=1}^{N}$ and $\{\hat{\mathbf{p}}_{n}\}_{n=1}^{N}$ be the ground‑truth and reconstructed frequency‑band patches defined in Section 2.2, and let $\mathcal{M}=\{n\mid m_{n}=1\}$ be the masked‑patch index set. The model is trained to minimise the mean Smooth‑L1 discrepancy over only the masked patches  [33, 34]:
@@
This objective 14 compels the network to recover the multiscale, frequency‑band information concealed by the mask.
@@ 3.1 Datasets and Training Settings
Leveraging the PhysioWave architecture, two large-scale, modality-specific foundation models, PhysioWave-ecg and PhysioWave-emg, are pretrained on the most extensive open-access corpora currently available for their respective signal types (see Tables 15 and 16). PhysioWave-ecg is trained on approximately 182 GB of twelve-lead ECG recordings, while PhysioWave-emg utilizes about 823 GB of EMG data. For each modality, we provide three parameter configurations: Small (5M), Base (15M), and Large (37M). Both PhysioWave models share the same Transformer backbone; however, their learnable wavelet front-ends are modality-aware (see Tables 12 and 13 in Appendix C for full architectural and training details).
@@
Each recording is denoised with a modality-specific band-pass filter followed by a 50 Hz notch. Traces with fewer than 12 ECG leads or 16 EMG electrodes are zero-padded and then resampled to 500 Hz (ECG) or 2 kHz (EMG). See Appendix D and E for full details.
@@ 3.2 Downstream Experiment Results
Table 1: ECG rhythm classification results on three benchmark datasets.
Method (year) Params PTB-XL CPSC 2018 Chapman-Shaoxing
F1 AUROC F1 AUROC F1 AUROC
ECG-Chat (2024) 13B 55.9 94.1 80.1 95.7
MERL (2024) 11M 48.1 91.9 72.8 92.6 87.9
MaeFE (2023) 9M 64.7 88.6 71.6 94.5
OpenECG-SimCLR 11M 46.9 91.5 73.1 92.4 52.3 95.1
OpenECG-BYOL 11M 47.7 91.1 72.8 92.6 51.5 94.8
OpenECG-MAE 11M 48.1 90.9 74.5 93.2 50.8 94.2
Ours–Small 5M 65.8 92.7 71.6 95.5 52.1 96.4
Ours–Base 15M 64.5 93.4 72.5 95.9 53.8 97.2
Ours–Large 37M 66.7 94.6 73.1 96.1 54.8 98.3
Note: Pink indicates the best results, blue indicates the second-best results.
Table 2: Surface-EMG gesture recognition performance across three datasets.
Method (year) Params NinaPro DB5 EPN-612 UCI EMG
Acc. F1 Acc. F1 Acc. F1
Moment (2024) 385M 86.41 74.42 90.87 90.16 90.45 91.75
OTiS (2024) 45M 85.31 72.61 87.55 88.03 90.62 89.28
Ours–Small 5M 84.78 72.54 93.12 93.40 90.35 89.51
Ours–Base 15M 86.02 73.78 93.68 93.91 91.92 92.77
Ours–Large 37M 87.53 75.42 94.50 94.56 93.19 93.59
  • Note: Pink indicates the best results, blue indicates the second-best results.
@@
On MPDB, the same multimodal architecture—obtained by replacing the EEG branch with LaBraM ($5.8$ M parameters)—improves accuracy from $70.4\%$ (using LaBraM alone) to $74.9\%$ ($4.5\%$), as shown in Figure 5. This outperforms generic sequence models such as MMPNet and EEGNet by $9\%$ to $17\%$  [37, 39].
Figure 5: Multimodal classification performance.
@@ 4 Conclusion
5 Ethics Statement
This study exclusively utilizes publicly available, fully de-identified physiological signal datasets (EEG, EMG, and ECG) with no collection of new data or access to personally identifiable information. All datasets were obtained in compliance with their respective terms of use, and the original data collections were conducted under appropriate institutional review board (IRB) approvals as documented in their publications: NinaPro (Ethics Commission of Canton Valais), MIMIC-IV-ECG (Beth Israel Deaconess Medical Center and MIT IRBs), PTB-XL (PTB Institutional Ethics Committee), DEAP (Queen Mary University of London Ethics Committee), and MPDB (Tsinghua University Medical Ethics Committee). Our secondary analysis of these anonymized public datasets adheres to the NeurIPS Ethics Guidelines, with no re-identification attempts made.
@@ 6 Acknowledgment
This research was partly supported by the EU Horizon Europe project IntelliMan (grant agreement 101070136), the PNRR MUR project ECS00000033 ECOSISTER, the ETH Zürich’s Future Computing Laboratory funded by a donation from Huawei Technologies, and also the EU Horizon Europe project HAL4SDV (grant agreement 101139789). In addition, we gratefully acknowledge the computational resources made available by Cineca and CSCS. In particular, this work utilised the GPU-based high-performance computing infrastructures provided by Cineca (Italy) and CSCS (Switzerland), whose support and operational services were instrumental in enabling the large-scale simulations/data-analysis undertaken for this study. Without their provision of computing time, maintenance of hardware and software stack, and user support, the computational component of our research would not have been possible.
Appendix A ADDITIONAL EXPERIMENTS
A.1 Ablation Study
To validate the effectiveness of each proposed component in PhysioWave, we conduct comprehensive ablation experiments by systematically removing or replacing specific sub-modules. All experiments are performed using the pretrained small PhysioWave model with fine-tuning on the target datasets.
A.2 Component Analysis
We analyze three key components: Soft Gate (SG), Channel Aggregation Feed-Forward Network (CAFFN), and Adaptive Wavelet (AW). When AW is disabled, we use a fixed db6 wavelet decomposition as the baseline.
Table 4: Main ablation study results on EPN612 dataset showing individual and combined component contributions.
Components Performance
SG CAFFN AW Accuracy (%) F1-Score (%) Improvement
89.88 89.76 -
91.71 91.66 $\uparrow$1.83%
91.53 91.46 $\uparrow$1.65%
92.23 92.22 $\uparrow$2.35%
93.12 93.10 $\uparrow$3.24%
92.64 92.89 $\uparrow$2.76%
92.52 92.48 $\uparrow$2.64%
93.85 93.51 $\uparrow$3.97%
SG = Soft Gate, CAFFN = Channel Aggregation FFN, AW = Adaptive Wavelet. Pink/blue highlights indicate the best and second-best results.
The results demonstrate that Adaptive Wavelet contributes the most individually (+2.35%), highlighting the importance of adaptive wavelet selection over fixed approaches. The complete model achieves optimal performance (93.85% accuracy), showing synergistic effects when all components are combined.
A.3 Gate Mechanism Comparison
We compare different gating strategies to validate our soft gate design choice.
Table 5: Gate mechanism comparison on EPN612 dataset.
Gate Type Threshold Accuracy (%) F1-Score (%) Accuracy Drop
Hard Gate 0.3 92.07 92.05 $\downarrow$1.78%
Hard Gate 0.5 92.25 92.23 $\downarrow$1.60%
Hard Gate 0.7 92.17 92.19 $\downarrow$1.68%
Gumbel Gate 0.5 91.47 91.51 $\downarrow$2.38%
Soft Gate - 93.85 93.51 -
Note: Accuracy drops are compared against the soft gate baseline. Pink indicates the best results, blue indicates the second-best results.
Our soft gate mechanism significantly outperforms alternatives, with hard gates showing threshold sensitivity and Gumbel gates exhibiting the largest performance degradation due to stochastic sampling being less suitable for physiological signals.
A.4 Feed-Forward Network Variants
We evaluate different FFN architectures to demonstrate the effectiveness of our CrossScale CAFFN design.
Table 6: Feed-forward network variant comparison on EPN612 dataset.
FFN Variant Accuracy (%) F1-Score (%) Accuracy Drop
Standard FFN 91.95 91.93 $\downarrow$1.90%
CAFFN 92.72 92.69 $\downarrow$1.13%
CrossScale No Attention 93.29 93.31 $\downarrow$0.56%
CrossScale CAFFN 93.85 93.51 -
Note: Accuracy drops are compared against CrossScale CAFFN. Pink indicates the best results, blue indicates the second-best results.
The progressive improvements from Standard FFN to CrossScale CAFFN demonstrate the value of each enhancement: channel aggregation (+0.77%), cross-scale fusion (+1.34%), and attention mechanisms (+1.90%).
A.5 Wavelet Selection Analysis
We investigate the impact of different wavelet configurations on both EMG and ECG modalities to validate our adaptive wavelet selection approach.
Table 7: Wavelet selection ablation study on EPN612 EMG dataset.
Wavelet Configuration Accuracy (%) F1-Score (%) Improvement
Random Xavier (Baseline) 92.25 92.22 -
Single Wavelet
db6 92.13 92.11 $\downarrow$0.12%
coif4 92.13 92.14 $\downarrow$0.12%
sym6 92.93 92.93 $\uparrow$0.68%
Multi-Wavelet Combinations
db6+sym6+coif4 93.51 93.51 $\uparrow$1.26%
db4+db6+sym4 93.90 93.90 $\uparrow$1.65%
db8+coif4+bior4.4 92.90 92.89 $\uparrow$0.65%
Note: Pink indicates the best results, blue indicates the second-best results.
Table 8: Wavelet selection ablation study on Georgia ECG dataset.
Wavelet Configuration Accuracy (%) F1-Score (%) Improvement
Random Xavier (Baseline) 62.38 48.75 -
Single Wavelet
db4 62.70 48.94 $\uparrow$0.32%
demy 63.59 49.36 $\uparrow$1.21%
sym4 62.45 48.99 $\uparrow$0.07%
Multi-Wavelet Combinations
db6+coif4+sym8 64.38 50.65 $\uparrow$2.00%
db8+sym5+coif3+demy 64.20 50.46 $\uparrow$1.82%
sym4+sym5+db6+coif3+bior4.4 63.70 49.93 $\uparrow$1.32%
Note: Pink indicates the best results, blue indicates the second-best results.
The results reveal that single wavelets provide minimal improvements (<1%), while multi-wavelet combinations achieve substantial gains (1.65% for EMG, 2.00% for ECG). Notably, optimal combinations differ between modalities: EMG benefits from db4+db6+sym4 for capturing transient muscle activations, while ECG prefers db6+coif4+sym8 for complex cardiac morphologies. This validates our modality-aware adaptive wavelet selection strategy.
A.6 Modality-Specific Adaptation for EEG and PPG Signals
To demonstrate the generalizability of our PhysioWave framework beyond ECG and EMG signals, we conducted experiments on EEG and PPG modalities. These experiments validate our model’s ability to adapt to diverse biosignal characteristics without requiring modality-specific architectural changes.
A.7 EEG Sleep Stage Classification
A.7.1 Dataset and Setup
We evaluated PhysioWave on the Sleep-EDF dataset for 5-class sleep stage classification[40]. EEG signals present unique challenges including low signal-to-noise ratio, high channel count, and significant heterogeneity across subjects. Our PhysioWave-EEG configuration uses diverse wavelet bases (db4, db6, sym4, coif2, bior2.2) suited for EEG characteristics, with a transformer backbone (embed_dim=512, depth=8, num_heads=8).
For the enhanced version, we incorporated three key improvements:
  • Multi-Scale Temporal Enhancement: Added multi-scale convolutions to capture EEG’s diverse frequency bands (delta, theta, alpha, beta, gamma) and improve SNR
  • Channel Interaction Modeling: Enhanced spatial relationship modeling between EEG electrodes through cross-channel attention
  • Window Attention: Implemented efficient processing of long EEG sequences with linear complexity
A.7.2 Results and Analysis
Table 9: EEG sleep stage classification results on Sleep-EDF dataset. All baseline methods use pretrained models, while PhysioWave is trained from scratch.
Method Balanced Accuracy Cohen’s Kappa Weighted F1
BENDR (pretrained) 0.6655 0.6659 0.7507
BIOT (pretrained) 0.6622 0.6461 0.7415
LaBraM (pretrained) 0.6771 0.6710 0.7592
EEGPT (pretrained) 0.6917 0.6857 0.7654
PhysioWave (original) 0.6720 0.6676 0.7472
PhysioWave (enhanced) 0.7312 0.7206 0.7839
Note: Pink indicates best results, blue indicates second-best results.
As shown in Table 9, even the original PhysioWave architecture achieves competitive performance against EEG-specific foundation models. The enhanced version achieves substantial improvements: +5.7% in Balanced Accuracy, +5.1% in Cohen’s Kappa, and +2.4% in Weighted F1 compared to the best baseline (EEGPT)[5]. Notably, these gains are achieved through direct training from scratch, without the computational overhead of large-scale pretraining required by baseline methods.
The performance improvements demonstrate that our wavelet-based approach effectively captures the multi-scale temporal dynamics inherent in EEG signals. The enhanced attention mechanisms successfully model inter-channel dependencies crucial for sleep stage classification, where different brain regions exhibit coordinated activity patterns.
A.8 PPG-Based Activity Recognition and Heart Rate Estimation
A.8.1 Dataset and Architecture
We evaluated PhysioWave on the PPG-DaLiA dataset, which contains PPG signals collected during various physical activities[41]. Given the deployment constraints of wearable devices, we designed PhysioWave-PPG as a compact variant optimized for edge computing:
  • Reduced Model Depth: 4 transformer layers (vs. 8-12 in standard configurations)
  • Smaller Embedding Dimensions: 192 (vs. 256-512)
  • PPG-Optimized Wavelets: db4, coif2, bior2.2 selected for cardiovascular periodicity
A.8.2 Activity Recognition Results
Table 10: Activity recognition performance on PPG-DaLiA dataset.
Method Params Accuracy Precision Recall F1 Score AUROC
Pulse-PPG (pretrained) 28.5M 0.4234 0.4398 0.3663 0.3648 0.8246
PhysioWave-PPG 2.53M 0.6209 0.6015 0.6077 0.5952 0.9420
Note: Pink indicates best results.
A.8.3 Heart Rate Regression Results
Table 11: Heart rate regression performance comparison.
Method Params MAE MSE MAPE
Pulse-PPG (pretrained) 28.5M 3.705 59.81 4.84%
Deep PPG 8.5M 7.65
NAS-PPG 0.8M 6.02
PhysioWave-PPG 1.57M 5.22 65.62 5.83%
Note: Pink indicates best results, blue indicates second-best results.
A.8.4 Performance Analysis
[…]
  • +63.1% F1-score improvement: 0.5952 vs 0.3648
  • +14.2% AUROC improvement: 0.9420 vs 0.8246
  • +46.7% accuracy improvement: 0.6209 vs 0.4234
[…]
A.9 Key Insights
These additional experiments demonstrate three critical capabilities of our PhysioWave framework:
  1. 1.
    Modality Generalization: The framework successfully adapts to signals with vastly different characteristics, from complex multichannel EEG with low SNR to periodic, single-channel PPG with clear cardiovascular patterns.
  2. 2.
    Efficiency Without Pretraining: PhysioWave achieves competitive or superior performance compared to pre-trained foundation models while training from scratch, eliminating the computational burden of large-scale pre-training.
  3. 3.
    Scalable Architecture: The framework scales effectively from compact edge deployments (1.57M parameters for PPG) to more complex clinical applications (enhanced EEG model), maintaining strong performance across the spectrum.
The success in the EEG and PPG modalities, combined with our primary ECG and EMG results, validates PhysioWave as a truly universal biosignal processing framework. The wavelet-transformer synergy provides the necessary flexibility to capture modality-specific patterns while maintaining a consistent architectural foundation.
Appendix B RELATED WORK
B.1 Self-supervised pre-training for Time Series